# MedRAT: Unpaired Medical Report Generation via Auxiliary Tasks

### v1

@@ Abstract
Generating medical reports for X-ray images is a challenging task, particularly in an unpaired scenario where paired image-report data is unavailable for training. To address this challenge, we propose a novel model that leverages the available information in two distinct datasets, one comprising reports and the other consisting of images. The core idea of our model revolves around the notion that combining auto-encoding report generation with multi-modal (report-image) alignment can offer a solution. However, the challenge persists regarding how to achieve this alignment when pair correspondence is absent. Our proposed solution involves the use of auxiliary tasks, particularly contrastive learning and classification, to position related images and reports in close proximity to each other. This approach differs from previous methods that rely on pre-processing steps using external information stored in a knowledge graph. Our model, named MedRAT, surpasses previous state-of-the-art methods, demonstrating the feasibility of generating comprehensive medical reports without the need for paired data or external tools.
@@ 1 Introduction
The automated generation of medical reports has the potential to improve the efficiency and accuracy of patient information documentation, ultimately leading to better patient care and resource savings. As a result, numerous studies have focused on this problem [3, 4, 15, 19, 37]. Although there are a few labeled image-report pair datasets available [6, 16], they are relatively small compared to datasets for natural image [23, 33, 34]. Additionally, these datasets might only be partially accessible to the public. For instance, the complete dataset of [14] could not be publicly released, resulting in a dataset containing solely images. This issue is attributed to privacy concerns, limited access to high-quality data, and the complexity of medical data analysis and labeling, which require expert knowledge. The shortage of labeled image-report pairs presents a significant challenge in this domain.
(a) Data                  (b) Unpaired Training             (c) Inference
@@
Figure 1: Unpaired medical report generation. (a) Two distinct and unpaired datasets are available for the task: chest X-ray images and chest X-ray reports. (b) We propose a model that addresses this challenge by learning to generate reports from reports, and embedding related images and reports close together in the embedding space. Notably, our model achieves this without access to image-report pairs during training. (c) By learning both tasks simultaneously, our model is able to generate detailed reports for X-ray images during inference.
@@
Unpaired methods for medical data may offer a solution to some of these concerns, in particular privacy issues and the shortage of paired data. This paper focuses on the task of medical report generation in an unpaired setting, which was first introduced in [27]. During training, two distinct datasets are available: one comprising chest X-ray images (e.g., [14]) and the other containing chest X-ray reports (e.g., [16]). Notably, the images and reports are not paired with each other. Each example in each dataset contains some domain information, typically related to the presence of specific pathologies. At inference, the model receives only an image and should generate a medical report. In [27], an auto-encoding pipeline is proposed to address this problem, which encodes shared information between text and images in a pre-constructed knowledge graph. Generating this hand-crafted graph requires expertise and exhibits a scalability weakness. Our approach eliminates the necessity for this graph.
@@
Our method presents two main ideas to generate medical reports for images, without paired image-report data. Firstly, it learns how to generate the same report given as an input, utilizing a language model and auto-encoding. Clearly, if we could only seamlessly replace the input report with an image during inference, our model would be able to generate reports from images. Hence, the second idea is to approach the domain gap by learning to represent the reports and the images in a shared space that captures similarities and dissimilarities between them, i.e. multi-modal alignment. Our novel architecture complements these goals by leveraging both global and local representations for the alignment and generation processes. Differently from [27], we do not rely on external tools, but instead solely use the provided data. If performed effectively, it would generate high-quality reports (Figure 1(c)). We elaborate on the two sub-tasks below.
@@
Although auto-encoding reports during training appears to be a straightforward approach, the limited diversity and size of the dataset can hinder generalization. To overcome this limitation, we propose a two-level approach. At the global level, the report representation focuses on capturing the overall information, in particular the presence of pathologies. At the local level, the representation exposes the model to detailed information such as the location, size, and relation to other organs or devices.
@@
Our method’s effectiveness is demonstrated through improvements in both language metrics and clinical metrics. For example, compared to the previous SoTA methods, it achieves a $14\%$ improvement in the BLEU-1 score (language efficacy) and a $10\%$ increase in recall (clinical efficacy) on the test dataset from [16]. In this experiment, we use the images from [14] for training; this dataset lacks public reports, hence cannot be used by paired methods.
@@ 2 Related Work
There is limited research on generating medical reports through unpaired methods. The closest related work is KGAE [27], which utilizes a pre-constructed knowledge graph to connect visual and textual modalities. However, graph design requires domain expertise. By training via auxiliary tasks we omit this requirement. We hereby elaborate on two related tasks.
@@ 4 Experimental Results
Datasets. We trained our model using images from the CheXpert dataset [14] and the training reports from the MIMIC-CXR dataset [16]. To evaluate the performance, we used test sets from two datasets: MIMIC-CXR and IU X-ray [6]. During training, we utilized the images only for the auxiliary tasks, while matching reports were only used for evaluation. Our experimental setup is similar to that of [27], with the exception that they also trained on reports from the IU X-ray dataset. We followed the same report preprocessing steps as [3, 27], where reports lacking a findings section were filtered out. Importantly, there were no paired samples between the CheXpert and MIMIC-CXR or IU X-ray datasets.
@@
Table 1: Quantitative evaluation, NLG metrics. Our results outperform those of [27] across all datasets & language generation metrics: BLEU (B), METEOR (M), ROUGE-L (R-L), RadGraph (R-G) & BERTScore (B-S). For the IU X-ray dataset, our model was not exposed to any data from the dataset during training, whereas [27] uses its reports.
@@ B-1 B-4 M R-L R-G B-S B-1 B-4 M R-L R-G B-S
Table 2: Quantitative evaluation, CE metrics. Our results outperform those of [27] in terms of clinical efficacy (CE) metrics, on the MIMIC-CXR dataset.
@@
Evaluation metrics. We assess our model’s performance in two aspects: the quality of the generated language (NLG), and the quality of its clinical efficacy (CE). To evaluate NLG, we utilize the common n-gram based BLEU [32], METEOR [1], and ROUGE-L [22] metrics, which measure the similarity between the generated reports and the ground-truth. We also employ two recent evaluation metrics studied for this task in [39], RadGraph & BERTScore, which show high alignment with radiologists judgment. To evaluate CE, we employ the CheXpert [14] model to assign $14$ diagnosis classes related to thoracic diseases and support devices, and we measure precision, recall, and F1 score with respect to the ground-truth labels.
@@
Quantitative evaluation. Table 1 presents a comparison between our method and [27], the only related work that addresses the same task, using NLG metrics. Our approach outperforms [27] in all relevant metrics. Additionally, our method achieves the results for the IU X-ray dataset in a zero-shot fashion, meaning that it was not exposed to any data from this dataset during training, neither images nor reports. On the other hand, [27] uses its reports for training. We note that since [27] have not yet released their code and the new metrics (RadGraph & BERTScore) are not reported we could not compare the results. Our reports not only resemble the ground-truth more closely, but are also more accurate and informative in extracting clinical information (pathologies), as shown in Table 2. (As explained in [3], these metrics do not suit IU X-ray dataset, due to its labeling schema, thus Table 2 focuses on MIMIC-CXR).
@@
Qualitative evaluation. The similarities between our generated report and the ground-truth are depicted in Figure 3. It is noteworthy that while the only global information associated with the image is the presence of "Edema", our report includes additional details such as the type and severity of the edema. Figure 4 displays attention maps that highlight the decoder’s focus while predicting specific words (blended with the input image). The maps show that the model identifies local connections between the image and the report when predicting words. For instance, the model focuses on the heart region while predicting the word "heart". Although the model was not trained on image-report pairs, the attention maps demonstrate its ability to establish relevant connections. (It is impossible to make qualitative comparisons to [27] since the code has not yet been released.)
@@
Figure 4: Attention visualization. (a) presents the input and (b-d) show examples of attention maps generated by our model, where bright values represent high attention. These maps demonstrate where the model is focusing when predicting specific words. For example, when predicting "heart" (b), the model’s attention is primarily on the central area of the image surrounding the heart; when predicting a tube "tip" in (c), the attention is concentrated up around the trachea, slightly above the carina; when predicting "pleural" in (d), the attention is focused on the left pleural cavity area (right side of the image). Notably, this is achieved even though our model was not trained on patch-word alignment.
@@ 6 Conclusions
This paper introduces a novel method for generating X-ray reports in an unpaired manner, without paired images and reports available during training. The proposed approach involves auto-encoding reports while simultaneously learning representations through two auxiliary tasks, namely contrastive learning and classification. These tasks leverage available domain knowledge for each modality. We demonstrate the effectiveness of our approach on two different datasets. It outperforms the state-of-the-art unpaired methods for chest X-ray report generation and significantly narrows the performance gap with paired methods.

### v2

@@ Abstract
Medical report generation from X-ray images is a challenging task, particularly in an unpaired setting where paired image-report data is unavailable for training. To address this challenge, we propose a novel model that leverages the available information in two distinct datasets, one comprising reports and the other consisting of images. The core idea of our model revolves around the notion that combining auto-encoding report generation with multi-modal (report-image) alignment can offer a solution. However, the challenge persists regarding how to achieve this alignment when pair correspondence is absent. Our proposed solution involves the use of auxiliary tasks, particularly contrastive learning and classification, to position related images and reports in close proximity to each other. This approach differs from previous methods that rely on pre-processing steps, such as using external information stored in a knowledge graph. Our model, named MedRAT, surpasses previous state-of-the-art methods, demonstrating the feasibility of generating comprehensive medical reports without the need for paired data or external tools.
@@ 1 Introduction
The automated generation of medical reports has the potential to improve the efficiency and accuracy of patient information documentation, ultimately leading to better patient care and resource savings. As a result, numerous studies have focused on this problem [3, 4, 17, 21, 40]. Although there are a few labeled image-report pair datasets available [7, 18], they are relatively small compared to datasets for natural image [25, 36, 37]. This is attributed to privacy concerns, limited access to high-quality data, and the complexity of medical data analysis and labeling, which require expert knowledge.
(a) Unpaired Training                               (b) Inference
@@
Figure 1: Unpaired medical report generation. (a) We propose a model that addresses the challenge of unpaired images and reports by learning to generate reports from reports, and embedding related images and reports close together in the embedding space. Notably, our model achieves this without access to image-report pairs during training. (b) By learning both tasks simultaneously, our model is able to generate detailed reports for X-ray images during inference.
@@
This paper focuses on the task of medical report generation in an unpaired setting, which addresses privacy issues and the shortage of paired data. During training, two distinct datasets are available: one comprising chest X-ray images (e.g., [16]) and the other containing chest X-ray reports (e.g., [18]). Notably, the images and reports are not paired with each other. Each example in each dataset contains some domain information, typically related to the presence of specific pathologies. At inference, the model receives only an image and should generate a medical report. The method of [29] relies on a pre-constructed knowledge graph, whereas the recent method suggested in [14] uses cycle-consistency and pseudo-reports. Our approach eliminates the necessity for designing the graph or pseudo-reports.
@@
Our method presents two main ideas to generate medical reports for images, without paired image-report data. Firstly, it learns how to generate the same report given as an input, utilizing a language model and auto-encoding. Clearly, if we could only seamlessly replace the input report with an image during inference, our model would be able to generate reports from images. Hence, the second idea is to approach the domain gap by learning to represent the reports and the images in a shared space that captures similarities and dissimilarities between them, i.e. multi-modal alignment. Our novel architecture complements these goals by leveraging both global and local representations for the alignment and generation processes. Differently from [14, 29], we do not rely on designing additional components (pseudo-reports or knowledge graphs), but instead solely use the provided data. If performed effectively, it would generate high-quality reports (Figure 1(b)). We elaborate on the two sub-tasks below.
@@
Although auto-encoding reports during training appears to be a straightforward approach, the limited diversity and size of the dataset can hinder generalization. To overcome this limitation, we propose a two-level approach. At the global level, the report representation focuses on capturing the overall information, in particular the presence of pathologies. At the local level, the representation exposes the model to detailed information such as the location, size, and relation to other organs or devices [6].
@@
Our method’s effectiveness is demonstrated through improvements in both language metrics and clinical metrics. For example, our method achieves a ${}10\%$ increase in clinical efficacy (recall) on the test dataset from [18] compared to [14, 29]. In this experiment, we use the images from [16] for training; this dataset lacks public reports, hence cannot be used by paired methods.
@@ 2 Related Work
There is limited research on generating medical reports through unpaired methods. Two closest related works are KGAE [29] and the recently published MedCycle [14]. KGAE utilizes a pre-constructed knowledge graph to connect visual and textual modalities. However, graph design requires domain expertise. MedCycle, instead, relies on cycle-consistency and pseudo-reports to bridge the two modalities. Consequently, the design of the pseudo-reports might significantly affect the results. In addition, mapping one modality to the other via cycle may still preserve hidden source information, just as a clue to map it back [34]. By training via auxiliary tasks in a shared space, we bypass these requirements. We hereby elaborate on two related tasks.
@@ 4 Experimental Results
Datasets. Our model was trained using images from the CheXpert dataset [16] and the training reports from the MIMIC-CXR dataset [18]. To evaluate the performance, we used test sets from two datasets: MIMIC-CXR and IU X-ray [7]. During training, we utilized the images only for the auxiliary tasks, while matching reports were only used for evaluation. Our experimental setup is similar to that of [29], with the exception that they also trained on reports from the IU X-ray dataset. We followed the same report preprocessing steps as [3, 29], where reports lacking a findings section were filtered out. Importantly, there were no paired samples between the CheXpert and MIMIC-CXR or IU X-ray datasets.
@@
Table 1: Quantitative evaluation, NLG metrics. Our results outperform those of [29, 14] for most datasets and language generation metrics: BLEU (B), METEOR (M), ROUGE-L (R-L), RadGraph F1 (R-G), and BERTScore (B-S).
@@ B-1 B-4 M R-L R-G B-S B-1 B-4 M R-L R-G B-S
Table 2: Quantitative evaluation, CE metrics. Our results outperform those of [29, 14] in terms of clinical efficacy (CE) metrics, on the MIMIC-CXR dataset.
@@
Evaluation metrics. We assess our model’s performance in two aspects: the quality of the generated language (NLG), and the quality of its clinical efficacy (CE). To evaluate NLG, we utilize the common n-gram based BLEU [35], METEOR [1], and ROUGE-L [24] metrics, which measure the similarity between the generated reports and the ground-truth. We also employ two recent evaluation metrics studied for this task in [42], RadGraph F1 & BERTScore, which show high alignment with radiologists judgment. To evaluate CE, we employ the CheXpert [16] model to assign $14$ diagnosis classes related to thoracic diseases and support devices, and we measure precision, recall, and F1 score with respect to the ground-truth labels.
@@
Quantitative evaluation. Table 1 presents a comparison between our method and previous works using NLG metrics. On the large and significant MIMIC-CXR dataset, our approach outperforms previous works in all metrics. On the IU X-ray dataset, our approach outperforms [29] in all metrics and [14] in some metrics, including RadGraph & BERTScore, which align more closely with radiologist judgment than the other metrics. We note that unlike [29], our method performs in a zero-shot fashion on this dataset, meaning that it was not exposed to any data from this dataset during training, neither images nor reports. Our reports not only resemble the ground-truth more closely, but are also more accurate and informative in extracting clinical information (pathologies), as shown in Table 2. (As explained in [3], these metrics do not suit IU X-ray dataset, due to its labeling schema, thus Table 2 focuses on MIMIC-CXR).
@@
Qualitative evaluation. The similarities between our generated report and the ground-truth are depicted in Figure 3. It is noteworthy that while the only global information associated with the image is the presence of "Edema", our report includes additional details such as the type and severity of the edema. Figure 4 displays attention maps that highlight the decoder’s focus while predicting specific words (blended with the input image). The maps show that the model identifies local connections between the image and the report when predicting words. For instance, the model focuses on the heart region while predicting the word "heart". Although the model was not trained on image-report pairs, the attention maps demonstrate its ability to establish relevant connections.
@@
Figure 4: Attention visualization. (a) presents the input and (b-d) show examples of attention maps generated by our model, where bright values represent high attention. These maps demonstrate where the model is focusing when predicting specific words. For example, when predicting "heart" (b), the model’s attention is on the central area around the heart; for "tip" (c), it concentrates around the trachea above the carina; and for "pleural" (d), it focuses on the left pleural cavity area (right side of the image). Notably, this is achieved without training on patch-word alignment.
@@ B-1 B-4 M B-1 B-4 M
Potential Impact. Our method utilizes unpaired data, mitigating the risk of patient privacy leaks. In turn, the opportunities in employing unpaired data may encourage data owners to release more data. However, it’s important to acknowledge potential risks inherent in automated systems, as they might be susceptible to errors. Given the significant impact of misdiagnoses, we advocate for these systems to complement radiologists rather than replace them.
@@ 6 Conclusions
This paper introduces a novel method for generating X-ray reports in an unpaired setting, without paired images and reports available during training. The proposed approach involves auto-encoding reports while simultaneously learning representations through two auxiliary tasks, namely contrastive learning and classification. These tasks leverage available domain knowledge for each modality. We demonstrate the effectiveness of our approach on two different datasets. It outperforms the state-of-the-art unpaired methods for chest X-ray report generation and significantly narrows the performance gap with paired methods.
Acknowledgements. We gratefully acknowledge the support of the Israel Science Foundation 2329/22.